# Polynomial Log-Marginals and Tweedie’s Formula

### v2

@@ 1 Empirical Bayes and compound decision
One of the main attractions of the empirical Bayes framework for the compound decision problem is that it can result in risk reduction without solving the deconvolution problem of estimating the common conditional distribution $F$. Rather, one could work with the observable marginal density $m(y)$, and use either parametric or nonparametric models to estimate it, and use the estimated $\hat{m}(y)$ as a plug-in in the empirical Bayes approach. This observable-only route computes estimators directly from $m(y)$, avoiding explicit priors and high-dimensional integration. For example, assuming $\mu_{i}$’s to be normally distributed with zero mean and unknown variance $\sigma_{0}^{2}$, leads to a Gaussian marginal for the $Y_{i}$’s. Using a method of moments estimator for the $\sigma_{0}^{2}$ leads to the famous, classical James–Stein estimator $\delta(y)=(1-(n-2)/\sum_{i=1}^{n}y_{i}^{2})y$ (Stein, 1956). Difficulty in choosing a parametric model for the marginal, leads the way to more flexible nonparametric models such as Kiefer & Wolfowitz (1956), which is easily seen to be convex and can benefit hugely from recent advances in convex optimization tools, as shown by Koenker & Mizera (2014). Yet another approach is proposed by Efron (Efron, 2008, 2012, 2011), using ‘Lindsey’s method’, which consists of estimating the marginal distribution from a Poisson regression of an integer order $K$, based on Lindsey’s Method that models the histogram bin counts of the observed $\bf y$}-values as Poisson random variables. Efron (2011)’s formulation yields locally adaptive shrinkage based on the local shape of the marginal histogram, while making no parametric assumptions on the prior. This has inspired subsequent developments in the empirical Bayes/compound decision problem by nonparametric density estimates of the marginal (e.g., Simon & Simon, 2013; Wager, 2014). As emphasized by Efron (2014), empirical Bayes admits two complementary strategies: f-modeling, which models the marginal density $m(\cdot)$ on the observation ($y$) scale, and g-modeling, which models the prior $F$ on the parameter ($\mu$) scale. The former have been relied upon heavily in applications, especially after Robbins (1956), while the latter have been predominant in the theoretical EB literature (Laird, 1978; Jiang & Zhang, 2009). Beyond EB, the Kiefer–Wolfowitz NPMLE can be viewed directly as a compound-decision rule; “no oracle needed” results due to (Ritov, 2024) show its Tweedie plug-in is minimax without reducing to an oracle Bayes benchmark. In high-dimensional prediction problems, it is also appealing to bypass the challenges of prior specification on a large parameter vector and instead work with the observable one-dimensional marginal $m(y)$ via Tweedie’s formula. Such a ‘predictive Bayes’ viewpoint leverages $m(y)$ directly for calibrated shrinkage and prediction without specifying a high-dimensional prior (padilla2018deconvolution). In a recent discussion, efron2020pea contrasts prediction with attribution, arguing that prediction is inherently easier and more forgiving than attribution, and is often ‘anti-parsimonious’ and ‘a-probabilistic,’ with theoretical efficiency replaced by empirical methods. He also shows that a two-step hierarchical approach—using Tweedie’s formula for the local false discovery rate in wide data settings (large $p$, small $n$) can drastically reduce dimensionality while avoiding explicit prior modeling and remaining transparent about the observable target. This motivates our focus on which functional forms of $m(\cdot)$ can arise from a bona fide Bayes model, and when the observable-only route, though appealing, might not always satisfy essential Bayes constraints.
@@ 1.1 Tweedie’s formula
Choosing $F$ to minimize $\mathrm{SURE}(F)$ provides a data-driven approximation to the oracle Bayes rule. Ghosh et al. (2025) further provide a unified ‘$g$-modeling’ (Efron, 2014) view showing that minimizing Stein’s unbiased risk estimate is equivalent (up to a constant) to Hyvárinen’s score matching (Hyvärinen & Dayan, 2005) for learning the prior, and establishes near-parametric convergence rates for regret/Fisher divergence with oracle inequalities under misspecification. For a comprehensive discussion of empirical Bayes and the Tweedie–Eddington formula, we refer readers to the lecture notes by Ignatiadis and Sen (ignatiadis_sen_empirical_bayes_2025).

### v3

@@ 1 Empirical Bayes and compound decision
One of the main attractions of the empirical Bayes framework for the compound decision problem is that it can result in risk reduction without solving the deconvolution problem of estimating the common conditional distribution $F$. Rather, one could work with the observable marginal density $m(y)$, and use either parametric or nonparametric models to estimate it, and use the estimated $\hat{m}(y)$ as a plug-in in the empirical Bayes approach. This observable-only route computes estimators directly from $m(y)$, avoiding explicit priors and high-dimensional integration. For example, assuming $\mu_{i}$’s to be normally distributed with zero mean and unknown variance $\sigma_{0}^{2}$, leads to a Gaussian marginal for the $Y_{i}$’s. Using a method of moments estimator for the $\sigma_{0}^{2}$ leads to the famous, classical James–Stein estimator $\delta(y)=(1-(n-2)/\sum_{i=1}^{n}y_{i}^{2})y$ (Stein, 1956). Difficulty in choosing a parametric model for the marginal, leads the way to more flexible nonparametric models such as Kiefer & Wolfowitz (1956), which is easily seen to be convex and can benefit hugely from recent advances in convex optimization tools, as shown by Koenker & Mizera (2014). Yet another approach is proposed by Efron (Efron, 2008, 2012, 2011), using ‘Lindsey’s method’, which consists of estimating the marginal distribution from a Poisson regression of an integer order $K$, based on Lindsey’s Method that models the histogram bin counts of the observed $\bf y$}-values as Poisson random variables. Efron (2011)’s formulation yields locally adaptive shrinkage based on the local shape of the marginal histogram, while making no parametric assumptions on the prior. This has inspired subsequent developments in the empirical Bayes/compound decision problem by nonparametric density estimates of the marginal (e.g., Simon & Simon, 2013; Wager, 2014). As emphasized by Efron (2014), empirical Bayes admits two complementary strategies: f-modeling, which models the marginal density $m(\cdot)$ on the observation ($y$) scale, and g-modeling, which models the prior $F$ on the parameter ($\mu$) scale. The former have been relied upon heavily in applications, especially after Robbins (1956), while the latter have been predominant in the theoretical EB literature (Laird, 1978; Jiang & Zhang, 2009). Beyond EB, the Kiefer–Wolfowitz NPMLE can be viewed directly as a compound-decision rule; “no oracle needed” results due to (Ritov, 2024) show its Tweedie plug-in is minimax without reducing to an oracle Bayes benchmark. In high-dimensional prediction problems, it is also appealing to bypass the challenges of prior specification on a large parameter vector and instead work with the observable one-dimensional marginal $m(y)$ via Tweedie’s formula. Such a ‘predictive Bayes’ viewpoint leverages $m(y)$ directly for calibrated shrinkage and prediction without specifying a high-dimensional prior (Madrid-Padilla et al., 2018). In a recent discussion, Efron (2020) contrasts prediction with attribution, arguing that prediction is inherently easier and more forgiving than attribution, and is often ‘anti-parsimonious’ and ‘a-probabilistic,’ with theoretical efficiency replaced by empirical methods. He also shows that a two-step hierarchical approach—using Tweedie’s formula for the local false discovery rate in wide data settings (large $p$, small $n$) can drastically reduce dimensionality while avoiding explicit prior modeling and remaining transparent about the observable target. This motivates our focus on which functional forms of $m(\cdot)$ can arise from a bona fide Bayes model, and when the observable-only route, though appealing, might not always satisfy essential Bayes constraints.
@@ 1.1 Tweedie’s formula
Choosing $F$ to minimize $\mathrm{SURE}(F)$ provides a data-driven approximation to the oracle Bayes rule. Ghosh et al. (2025) further provide a unified ‘$g$-modeling’ (Efron, 2014) view showing that minimizing Stein’s unbiased risk estimate is equivalent (up to a constant) to Hyvárinen’s score matching (Hyvärinen & Dayan, 2005) for learning the prior, and establishes near-parametric convergence rates for regret/Fisher divergence with oracle inequalities under misspecification. For a comprehensive discussion of empirical Bayes and the Tweedie–Eddington formula, we refer readers to the lecture notes by Ignatiadis and Sen (Ignatiadis & Sen, 2025).